OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

-2y+4

OpenStudy (anonymous):

@speedygonzales can you help me?

OpenStudy (anonymous):

-x - 2y = -4 3x + y = 12 x=-2y+4 3(-2y+4)+y = 12 -6y+12+y=12 -5y = 0 y = 0 -x-2(0) = -4 -x = -4 x = 4 (4,0)

OpenStudy (anonymous):

ok, thanks